OpenStudy (anonymous):

find the measure of angle descibed. the angle's measure is 20 degrees more then thatof complement

OpenStudy (anonymous):

hello help please

Directrix (directrix):

Let x = the measure of one of the complementary angles Then 90 - x = the other x = (90-x) + 20 x = 90 - x + 20 x = 110 - x 2x = 110 x = 55 90 - x = 35 Answer --> 55

Directrix (directrix):

Question?